OpenStudy (anonymous):

The slope of the line whose equation is y - 3 = 0 is

OpenStudy (anonymous):

A. 0 B. 3 C. No slope

OpenStudy (amistre64):

do you know the slope intercept form of a line?

OpenStudy (anonymous):

y = mx + b

OpenStudy (amistre64):

good, and you are aware that m refers to the slope right?

OpenStudy (anonymous):

Yes.

OpenStudy (amistre64):

y = mx + b y = 0x + b y = b what is the slope of a line that has no x parts?

OpenStudy (anonymous):

Undefined?

OpenStudy (amistre64):

y = 0x + b ^ this is defined

OpenStudy (anonymous):

So 0?

OpenStudy (amistre64):

Zero works for me

OpenStudy (amistre64):

something is undefined when it is not even zero .... \(\not 0\)

OpenStudy (anonymous):

Okay. I get those two confused a lot. So thanks.

OpenStudy (amistre64):

:) good luck

OpenStudy (anonymous):

Thanks!